OpenStudy (anonymous):

1. A book 500 pages (500 sheets of paper) long is 2.25 inches thick. If each cover is 0.125 inches thick, how thick is each sheet of paper? 2.Jack buys a magic bean plant that is 6 inches tall and grows several inches a day. After 48 days the bean plant is 198 inches tall. How many inches does the bean plant grow each day? 3.Dan takes a test worth 100 points. Each question is worth the same number of points. Dan misses 5 questions and scores 80 on the test. How many points is each question worth? 4.Scott has a jar of quarters and dimes. If there is $4.15 in the jar, and there are 4 dimes, how many quarters are in the jar? 5.The volume of a rectangular prism is 2,100 cm3 with length 25 cm and width 12 cm. What is the height, in cm?

OpenStudy (anonymous):

just answers please, after i answer it it will show me a vidoe on step by step solving of that equation and ones like it. thank you

OpenStudy (ganpat):

1. Cover is what that covers the book, more like a cardboard.. so basically there are two cover on the front and at the back.. Soo, total thickness per page = [ total thickness - 2(cover thickness) / total pages]

OpenStudy (ganpat):

5. Volume of rectangle = l *b * h 2100 = 25 * 12 * h ... find h !!

OpenStudy (anonymous):

number 1. 0.4 number 2. 6+48x=198 it is already 6 inches tall. 48 * x is the length by days. total would be the length number 3. i am thinking that the answer is 4 because if he missed 5 questons and each question is worth 4 points 5x4=20 and 20 from 100 is 80

OpenStudy (anonymous):

that u so much @bd333140

OpenStudy (anonymous):

your welcome number 4. 25x + 40= 415 number 5. 7